xe^y dy + x^2+1/y dx= 0
1
Expert's answer
2021-10-26T17:21:20-0400
"xe^y dy + \dfrac{x^2+1}{y} dx= 0"

"ye^ydy=-\dfrac{x^2+1}{x}dx"

Integrate both sides


"\int ye^ydy=-\int\dfrac{x^2+1}{x}dx"

"\int ye^ydy=ye^y-\int e^ydy=ye^y-e^y+C_1"

"ye^y-e^y=-\dfrac{x^2}{2}-\ln x+C"